### v2

# FormCoach: Lift Smarter, Not Harder

Abstract
Good form is the difference between strength and strain, yet for the fast-growing community of at-home fitness enthusiasts, expert feedback is often out of reach. FormCoach transforms a simple camera into an always-on, interactive AI training partner, capable of spotting subtle form errors and delivering tailored corrections in real time, leveraging vision-language models (VLMs). We showcase this capability through a web interface and benchmark state-of-the-art VLMs on a dataset of 1,700 expert-annotated user–reference video pairs spanning 22 strength and mobility exercises. To accelerate research in AI-driven coaching, we release both the dataset and an automated, rubric-based evaluation pipeline, enabling standardized comparison across models. Our benchmarks reveal substantial gaps from human-level coaching, underscoring both the challenges and opportunities in bringing nuanced, context-aware movement analysis into interactive AI systems. By framing form correction as a collaborative, creative process between human and machine, FormCoach opens a new frontier in embodied AI.
1 Introduction
The only thing worse than not lifting weights is lifting them with bad form Jones et al. (1994). Poor techniques not only diminishes training effectiveness, it can also lead to serious injury Aasa et al. (2016). Yet, most individuals lack continuous access to a personal coach who can spot and correct errors on the spot Tsiouris et al. (2020); Braga-Pereira et al. (2024). Without timely, precise feedback, at-home fitness enthusiasts risk reinforcing bad habits, limiting progress, and an increasing likelihood of harm Graves et al. (2013). We introduce FormCoach, an interactive AI form coach designed to address exactly this challenge. FormCoach transform a simple camera, whether in a smartphone, tablet, or smart mirror, into an always-available form coach. Users can specify personalized goals such as "focus on my knee alignment," perform an exercise, and receive immediate, concise feedback. By comparing the user’s movement in real time against an expert reference video, FormCoach delivers actionable corrections tailored to the individual.
Previous AI trainer systems Fieraru et al. (2021); Chen and Yang (2020); Liu et al. (2024); Dittakavi et al. (2022) largely focused on pose estimation and joint angle differences, overlooking richer contextual cues such as body-equipment interaction and ground contact. Motion-language models have made significant progress in motion captioning, generation, and editing Athanasiou et al. (2024); Zhang et al. (2024); Chen et al. (2024a). Yet their progress is constrained by the scarcity of high-quality motion-language datasets, and they rarely address the problem of delivering concise, context-aware comparative feedback. Recent advances in vision-language models (VLMs) have unlocked impressive capabilities in video captioning, question answering, and multimodal reasoning OpenAI (2025a); Boqiang Zhang (2025); Chen et al. (2024b); Yang et al. (2024). However, they have rarely been tested on nuanced, comparative analyses of human motion Cho et al. (2025). FormCoach bridges this gap by evaluating state-of-the-art VLMs on a curated dataset of user-reference exercise pairs, probing their ability to identify and articulate form differences with clarity and precision. In doing so, FormCoach reframes AI as an interactive, creative partner, one that collaborates with users to elevate the at-home fitness experience.
Figure 1: FormCoach pipeline
2 AI-powered form coaching
FormCoach transforms a solitary home workout into an interactive coaching session powered by vision–language models that sees, understands, and responds like a human trainer. The FormCoach pipeline is illustrated in Figure 1.
Figure 2: Prototype web interface
Setup The user positions a camera (be it a webcam, smartphone camera, or smart mirror) to capture their workout space. From an extensive, curated library of exercises, the user selects an activity, optionally typing in personal goals or coaching preferences such as “focus on my knee alignment” or “help me improve my depth in squats.” These inputs are incorporated directly into the VLM prompt, shaping the system’s analysis and ensuring feedback is tailored to the user’s priorities. Once configured, FormCoach displays an expert demonstration video alongside the user’s real-time reflection, creating an immediate, side-by-side visual comparison (see Figure 2).
Perform As the user begins, FormCoach continuously tracks their body movements, synchronizing frames from the live feed with frames from the expert reference. These frames along with the user’s stated preferences, are analyzed by carefully prompted VLMs to detect form deviations.
Feedback When discrepancies appear, FormCoach responds with targeted, actionable guidance. For instance, if a user’s back rounds during a squat, the system might say, "Keep your back straight." Feedback is concise, on-screen, and optionally delivered via text-to-speech for a seamless, hands-free experience that doesn’t disrupt the workout flow.
3 Benchmark
3.1 Dataset
Source collection Our evaluation dataset uses Fit3D’s Fieraru et al. (2021) multi-view exercise recordings. We segmented each sequence into single-rep clips (e.g. one squat). We pair identical exercises to form 1010k reference-user video pairs, each aligned to a fixed-window (mean $\sim$3.8 s) and standardized to $224\times 244$ at $30$ FPS. We then uniformly sampled $1,700$ pairs as our evaluation dataset.
Exercise coverage The dataset covers $22$ exercises, summarized in Table 3, covering upper body lifts (e.g. dumbbell lift overhead), lower body movements (e.g. squats, lunges) and full body exercises (e.g. mule kick, burpee). The dataset is class-balanced, challenging models to perform consistently across exercise types rather than exploiting class frequency.
Ground-truth instructions Expert annotators are given the user and reference videos side by side on the gradio interface Abid et al. (2019) and asked to provide concise (<<\!15 words), actionable corrections comparing the paired movements (e.g., “Keep your back straight throughout the squat.”). These annotations define the FormCoach evaluation set of 1,700 pairs and are included in our release.
Figure 3: Example from FormCoach evaluation dataset showing the reference (top) and user (bottom) performing the same squatting exercise. Below, we present the expert feedback and sample VLM-generated feedbacks. We highlight the correct feedback in green, hallucinated feedback in yellow, and incorrect/unclear feedback in red.
Expert
"Spread your feet wider and push your knees outward as you squat down."
GPT-4.1 OpenAI (2025a)
Widen your stance and push your knees out as you squat to improve depth and stability.”
QwenVL2.5-7B Yang et al. (2024)
"Keep your back straight and core engaged."
VideoLLaMA3 Boqiang Zhang (2025)
“Keep your chest up and back straight.
Gemini-2.0- Hassabis et al. (2024)
“Keep your chest up and ensure your hips break the parallel line to maximize glute activation.”
3.2 Task Description
We experiment with various VLMs to generate form feedback for exercise video pairs from our dataset, as seen in Figure 3. These models included both API-based models OpenAI (2025a); Anthropic (2025); Hassabis et al. (2024), and open-source alternatives Boqiang Zhang (2025); Chen et al. (2024b); GLM et al. (2024); Yang et al. (2024). Each query consists of a system prompt and a user-reference exercise video pair. The models are instructed to generate targeted feedback focusing on form corrections within a limit of 15 words. The experimentation encompassed 1,700 exercise pairs, representing a diverse set of exercise types to ensure robust coverage. We ran the API-based models on CPUs and the open-source models on a single NVIDIA A40 GPU with 48 GB memory.
3.3 Evaluation
The quality of feedback generated by each model was compared against ground-truth annotations. We leverage GPT-4.1 OpenAI (2025a) for evaluation. GPT is presented with the annotated feedback and the model’s feedback (with relevant context) and asked to rate the model’s feedback on several dimensions. All GPT evaluations are done via the OpenAI API with a deterministic setting (temperature 0) to ensure consistency. The metrics described below:
Table 1: LLM-based evaluation metrics
Metric Values Description
Accuracy 5 Fully correct (pinpoints actual issue)
4 Mostly correct (minor details missing)
3 Partially correct (some missing or incorrect)
2 Mostly incorrect (mostly missing or incorrect)
1 Mostly correct (misses the issue)
Actionability 5 Very clear, specific, and actionable
4 Mostly clear, specific, and actionable
3 Somewhat clear, specific, and actionable
2 Mostly unclear or unsafe
1 Vague or unsafe
Hallucination Yes/No Whether the model flagged a form error that wasn’t actually present.
4 Results and discussions
Table 2: LLM-based evaluation results for different vision-language models.
Model Accuracy (%) $\uparrow$ Actionability (%) $\uparrow$ Hallucination (%) $\downarrow$
GPT-4.1 OpenAI (2025a) 58.2 94.4 74.26
GPT-o4-mini OpenAI (2025b) 56.2 86.4 93.99
Claude Opus 4 Anthropic (2025) 49.6 84.6 82.79
QwenVL2.5-7B Yang et al. (2024) 46.6 79.8 94.39
QwenVL2.5-3B Yang et al. (2024) 45.8 87.4 96.79
InternVL2.5-8B Chen et al. (2024b) 46.4 82.4 93.60
InternVL2.5-4B Chen et al. (2024b) 44.6 87.0 95.60
InternVL2.5-2B Chen et al. (2024b) 42.6 83.4 94.16
GLM4V-9B GLM et al. (2024) 45.0 85.6 94.62
VideoLLaMA3 Boqiang Zhang (2025) 44.8 69.0 75.98
Gemini-2.0-Flash Hassabis et al. (2024) 43.8 91.2 97.3
4.1 Performance gaps and opportunities
While the evaluated models show encouraging potential, they still fall short of expert human feedback. GPT-4.1 OpenAI (2025a), top performer, achieved an accuracy score of 58.2 and an actionability score of 94.4, meaning it consistently delivers clear instructions but correctly identifies only a little over half of actual form issues. The gap highlights room for improvement in precisely identifying and articulating form discrepancies. Other models lagged further behind, with lower accuracy and higher hallucination rates, sometimes inventing corrections unrelated to the reference form or missing mistakes entirely.
These shortcomings are not surprising given the inherent challenges of video-only analysis. Relying solely on visual input makes the system vulnerable to occlusions (e.g., obscured joints or equipment) and viewpoint limitations, which hinder detection of fine-grained posture errors. Even when movement differences exist, they can appear ambiguous or indistinguishable from certain camera angles.
These limitations indicate clear directions for enhancement. Multi-modal augmentation, such as combining video with 3D joint position estimationCao et al. (2019); Toshev and Szegedy (2014), could add spatial depth and enhance the model’s sensitivity to subtle kinematic differences. Furthermore, integrating wearable sensors Zhou et al. (2016), such as inertial measurement units (IMUs) Seuter et al. (2016), can provide precise and viewpoint-independent motion data. These enhancements would substantially reduce ambiguity, boost detection accuracy, and close the gap between AI and expert coaching, unlocking a new level of reliability and trust in AI-powered fitness guidance.
4.2 Exercise-wise breakdown
We computed the mean accuracy score for each exercise across all evaluated VLMs, highlighting the three highest- and three lowest-performing cases in Figure 4. The top performers (barbell shrug, hands-up rotate, and opposite arms up-and-down warmup) are predominantly isolated, upper-body movements. These exercises are visually straightforward, involving a single limb group with minimal self-occlusion, clear lines of sight, and relatively simple motion patterns. This suggests that VLMs excel when analyzing movements with distinct, easily trackable features and limited inter-limb interference.
In contrast, the lowest-performing exercises (burpees, barbell pull-up, and mule kick) are complex, full-body activities characterized by multiple movement phases, rapid posture transitions, and simultaneous coordination of several limb groups. These dynamics introduce frequent occlusion, abrupt viewpoint changes, and more complex spatial relationships between joints, all of which challenge current VLM architectures. The pronounced drop in accuracy for these tasks underscores an important limitation: today’s VLMs are not yet adept at reliably parsing the fluid, multi-stage mechanics of compound movements.
Figure 4: Accuracy of VLM-generated Feedback Across Exercises: Top 3 vs. Bottom 3
5 The future of at-home fitness coaching
From Feedback to Collaboration: A key opportunity lies in shifting from one-way feedback to two-way collaboration. In its current form, FormCoach primarily tells the user what to fix, but true coaching is often a dialogue. We envision future AI fitness partners that not only give instructions, but also listen to the user’s concerns, goals, and progress Garbett et al. (2021); Luo et al. (2021). By integrating wearable sensing for real-time state detection with conversational AI, such systems could adapt feedback dynamically, offering not just corrections, but strategic guidance tailored to the user’s intent and physical condition.
Augmented Reality and Embodiment: Advances in AR hardware and software are rapidly opening the door to immersive, embodied coaching experiences Ghosh et al. (2022); Pedagogue (2025). In an AR-enhanced FormCoach, lightweight smart glasses could project a virtual instructor directly into the user’s environment, visually overlaying highlights where their form diverges from the reference. Spatial audio could make feedback even more intuitive: imagine hearing your coach’s voice coming from the direction of a misaligned limb. This merging of spatial computing and AI analysis creates the potential for fully embodied, in-the-moment guidance that feels as natural as training with a human partner.
Accessible Fitness Coaching With AI coaches, expert guidance could become accessible to anyone with a camera, democratizing high-quality fitness training across geographies and income levels Magazine (2025). FormCoach envisions a future where every home gym includes a built-in AI trainer Engineer (2025), much like AI assistants are now standard in smart homes Brown (2019); Cottier (2024). The challenge will be to ensure that as these systems become more pervasive, their feedback remains accurate, safe, and personalized—even as users push their physical limits.
6 Conclusion
We presented FormCoach, an interactive AI form coach capable of delivering targeted, actionable exercise feedback in real time. Alongside the system, we developed a curated dataset of expert-annotated user–reference exercise pairs and used it to systematically evaluate state-of-the-art vision–language models (VLMs) on nuanced, comparative human motion analysis. Our results demonstrate that VLMs can move beyond generic video understanding toward precise, context-aware guidance, positioning AI as an interactive, creative coaching partner. Our experiments show that while top-performing models such as GPT-4.1 exhibit strong actionability and partial accuracy, they still miss subtle form discrepancies and occasionally hallucinate nonexistent issues. These findings highlight both the promise and the limitations of current VLMs in high-stakes, precision-oriented applications. Overall, FormCoach opens the door to a new class of AI systems that not only see but also engage with users, guiding them through skill acquisition in a collaborative, human-like manner. As multimodal AI advances and incorporates richer sensing, the vision of universally accessible, trustworthy AI coaches moves closer to reality.
Appendix A Technical Appendices and Supplementary Material
A.1 Dataset
The dataset includes 22 common strength and mobility exercises as listed in Table 3. The dataset is hosted on Huggingface and can be accessed Here.
Table 3: List of Exercises Included in the Dataset
No. Exercise Name
1 Dumbbell Lift Overhead
2 Dumbbell Lunge Backward
3 Hands Up Rotate
4 Barbell Pull Up
5 Clean and Press
6 Dumbbell Lateral Raise
7 Burpee
8 Up and Down Stretch
9 Dumbbell Chest Fly
10 Arms Raise Overhead
11 Standing Ab Twists
12 Opposite Arms Up and Down Warmup
13 Dumbbell Squat and Press
14 Lunge and Reach Hands Front
15 Mule Kick
16 Arms Circles
17 Squat
18 One Arm Row
19 Squat and Jump with Hands Up
20 Dumbbell High Pulls
21 Deadlift
22 Barbell Shrug
A.2 Ground Truth Instruction Annotations
Experts annotated ground truth instructions given the user and reference exercise pairs through the Gradio interface as illustrated in Figure  5.
Figure 5: Instruction Annotation Interface on Gradio
A.3 Prompts
Prompts for feedback generation and evaluations are shown in Figure 6, Figure 8, Figure 8, and Figure 9.
Figure 6: Prompt for feedback generation
Figure 7: Prompt for accuracy evaluation
Figure 8: Prompt for actionability evaluation
Figure 9: Prompt for hallucination evaluation