### v3

# FormCoach: Lift Smarter, Not Harder

Abstract
Good form is the difference between strength and strain, yet for the fast-growing community of at-home fitness enthusiasts, expert feedback is often out of reach. FormCoach transforms a simple camera into an always-on, interactive AI training partner, capable of spotting subtle form errors and delivering tailored corrections in real time, leveraging vision-language models (VLMs). We showcase this capability through a web interface and benchmark state-of-the-art VLMs on a dataset of 1,700 expert-annotated user–reference video pairs spanning 22 strength and mobility exercises. To accelerate research in AI-driven coaching, we release both the dataset and an automated, rubric-based evaluation pipeline, enabling standardized comparison across models. Our benchmarks reveal substantial gaps from human-level coaching, underscoring both the challenges and opportunities in bringing nuanced, context-aware movement analysis into interactive AI systems. By framing form correction as a collaborative, creative process between human and machine, FormCoach opens a new frontier in embodied AI.
1 Introduction
The only thing worse than not lifting weights is lifting them with bad form [1]. Poor techniques not only diminishes training effectiveness, it can also lead to serious injury [2]. Yet, most individuals lack continuous access to a personal coach who can spot and correct errors on the spot [3, 4]. Without timely, precise feedback, at-home fitness enthusiasts risk reinforcing bad habits, limiting progress, and an increasing likelihood of harm [5]. We introduce FormCoach, an interactive AI form coach designed to address exactly this challenge. FormCoach transform a simple camera, whether in a smartphone, tablet, or smart mirror, into an always-available form coach. Users can specify personalized goals such as "focus on my knee alignment," perform an exercise, and receive immediate, concise feedback. By comparing the user’s movement in real time against an expert reference video, FormCoach delivers actionable corrections tailored to the individual.
Previous AI trainer systems [6, 7, 8, 9] largely focused on pose estimation and joint angle differences, overlooking richer contextual cues such as body-equipment interaction and ground contact. Motion-language models have made significant progress in motion captioning, generation, and editing [10, 11, 12]. Yet their progress is constrained by the scarcity of high-quality motion-language datasets, and they rarely address the problem of delivering concise, context-aware comparative feedback. Recent advances in vision-language models (VLMs) have unlocked impressive capabilities in video captioning, question answering, and multimodal reasoning [13, 14, 15, 16]. However, they have rarely been tested on nuanced, comparative analyses of human motion [17]. FormCoach bridges this gap by evaluating state-of-the-art VLMs on a curated dataset of user-reference exercise pairs, probing their ability to identify and articulate form differences with clarity and precision. In doing so, FormCoach reframes AI as an interactive, creative partner, one that collaborates with users to elevate the at-home fitness experience.
Figure 1: FormCoach pipeline
2 AI-powered form coaching
FormCoach transforms a solitary home workout into an interactive coaching session powered by vision–language models that sees, understands, and responds like a human trainer. The FormCoach pipeline is illustrated in Figure 1.
Figure 2: Prototype web interface
Setup The user positions a camera (be it a webcam, smartphone camera, or smart mirror) to capture their workout space. From an extensive, curated library of exercises, the user selects an activity, optionally typing in personal goals or coaching preferences such as “focus on my knee alignment” or “help me improve my depth in squats.” These inputs are incorporated directly into the VLM prompt, shaping the system’s analysis and ensuring feedback is tailored to the user’s priorities. Once configured, FormCoach displays an expert demonstration video alongside the user’s real-time reflection, creating an immediate, side-by-side visual comparison (see Figure 2).
Perform As the user begins, FormCoach continuously tracks their body movements, synchronizing frames from the live feed with frames from the expert reference. These frames along with the user’s stated preferences, are analyzed by carefully prompted VLMs to detect form deviations.
Feedback When discrepancies appear, FormCoach responds with targeted, actionable guidance. For instance, if a user’s back rounds during a squat, the system might say, "Keep your back straight." Feedback is concise, on-screen, and optionally delivered via text-to-speech for a seamless, hands-free experience that doesn’t disrupt the workout flow.
3 Benchmark
3.1 Dataset
Source collection Our evaluation dataset uses Fit3D’s [6] multi-view exercise recordings. We segmented each sequence into single-rep clips (e.g. one squat). We pair identical exercises to form 1010k reference-user video pairs, each aligned to a fixed-window (mean $\sim$3.8 s) and standardized to $224\times 244$ at $30$ FPS. We then uniformly sampled $1,700$ pairs as our evaluation dataset.
Exercise coverage The dataset covers $22$ exercises, summarized in Table 3, covering upper body lifts (e.g. dumbbell lift overhead), lower body movements (e.g. squats, lunges) and full body exercises (e.g. mule kick, burpee). The dataset is class-balanced, challenging models to perform consistently across exercise types rather than exploiting class frequency.
Ground-truth instructions Expert annotators are given the user and reference videos side by side on the gradio interface [18] and asked to provide concise (<<\!15 words), actionable corrections comparing the paired movements (e.g., “Keep your back straight throughout the squat.”). These annotations define the FormCoach evaluation set of 1,700 pairs and are included in our release.
Figure 3: Example from FormCoach evaluation dataset showing the reference (top) and user (bottom) performing the same squatting exercise. Below, we present the expert feedback and sample VLM-generated feedbacks. We highlight the correct feedback in green, hallucinated feedback in yellow, and incorrect/unclear feedback in red.
Expert
"Spread your feet wider and push your knees outward as you squat down."
GPT‑4.1 [13]
Widen your stance and push your knees out as you squat to improve depth and stability.”
QwenVL2.5-7B [16]
"Keep your back straight and core engaged."
VideoLLaMA3 [14]
“Keep your chest up and back straight.
Gemini-2.0- [19]
“Keep your chest up and ensure your hips break the parallel line to maximize glute activation.”
3.2 Task Description
We experiment with various VLMs to generate form feedback for exercise video pairs from our dataset, as seen in Figure 3. These models included both API-based models [13, 20, 19], and open-source alternatives [14, 15, 21, 16]. Each query consists of a system prompt and a user-reference exercise video pair. The models are instructed to generate targeted feedback focusing on form corrections within a limit of 15 words. The experimentation encompassed 1,700 exercise pairs, representing a diverse set of exercise types to ensure robust coverage. We ran the API-based models on CPUs and the open-source models on a single NVIDIA A40 GPU with 48 GB memory.
3.3 Evaluation
The quality of feedback generated by each model was compared against ground-truth annotations. We leverage GPT-4.1 [13] for evaluation. GPT is presented with the annotated feedback and the model’s feedback (with relevant context) and asked to rate the model’s feedback on several dimensions. All GPT evaluations are done via the OpenAI API with a deterministic setting (temperature 0) to ensure consistency. The metrics described below:
Table 1: LLM-based evaluation metrics
Metric Values Description
Accuracy 5 Fully correct (pinpoints actual issue)
4 Mostly correct (minor details missing)
3 Partially correct (some missing or incorrect)
2 Mostly incorrect (mostly missing or incorrect)
1 Mostly correct (misses the issue)
Actionability 5 Very clear, specific, and actionable
4 Mostly clear, specific, and actionable
3 Somewhat clear, specific, and actionable
2 Mostly unclear or unsafe
1 Vague or unsafe
Hallucination Yes/No Whether the model flagged a form error that wasn’t actually present.
4 Results and discussions
Table 2: LLM-based evaluation results for different vision-language models.
Model Accuracy (%) $\uparrow$ Actionability (%) $\uparrow$ Hallucination (%) $\downarrow$
GPT-4.1 [13] 58.2 94.4 74.26
GPT-o4-mini [22] 56.2 86.4 93.99
Claude Opus 4 [20] 49.6 84.6 82.79
QwenVL2.5-7B [16] 46.6 79.8 94.39
QwenVL2.5-3B [16] 45.8 87.4 96.79
InternVL2.5-8B [15] 46.4 82.4 93.60
InternVL2.5-4B [15] 44.6 87.0 95.60
InternVL2.5-2B [15] 42.6 83.4 94.16
GLM4V-9B [21] 45.0 85.6 94.62
VideoLLaMA3 [14] 44.8 69.0 75.98
Gemini-2.0-Flash [19] 43.8 91.2 97.3
4.1 Performance gaps and opportunities
While the evaluated models show encouraging potential, they still fall short of expert human feedback. GPT-4.1 [13], top performer, achieved an accuracy score of 58.2 and an actionability score of 94.4, meaning it consistently delivers clear instructions but correctly identifies only a little over half of actual form issues. The gap highlights room for improvement in precisely identifying and articulating form discrepancies. Other models lagged further behind, with lower accuracy and higher hallucination rates, sometimes inventing corrections unrelated to the reference form or missing mistakes entirely.
These shortcomings are not surprising given the inherent challenges of video-only analysis. Relying solely on visual input makes the system vulnerable to occlusions (e.g., obscured joints or equipment) and viewpoint limitations, which hinder detection of fine-grained posture errors. Even when movement differences exist, they can appear ambiguous or indistinguishable from certain camera angles.
These limitations indicate clear directions for enhancement. Multi-modal augmentation, such as combining video with 3D joint position estimation[23, 24], could add spatial depth and enhance the model’s sensitivity to subtle kinematic differences. Furthermore, integrating wearable sensors [25], such as inertial measurement units (IMUs) [26], can provide precise and viewpoint-independent motion data. These enhancements would substantially reduce ambiguity, boost detection accuracy, and close the gap between AI and expert coaching, unlocking a new level of reliability and trust in AI-powered fitness guidance.
4.2 Exercise‑wise breakdown
We computed the mean accuracy score for each exercise across all evaluated VLMs, highlighting the three highest- and three lowest-performing cases in Figure 4. The top performers (barbell shrug, hands-up rotate, and opposite arms up-and-down warmup) are predominantly isolated, upper-body movements. These exercises are visually straightforward, involving a single limb group with minimal self-occlusion, clear lines of sight, and relatively simple motion patterns. This suggests that VLMs excel when analyzing movements with distinct, easily trackable features and limited inter-limb interference.
In contrast, the lowest-performing exercises (burpees, barbell pull-up, and mule kick) are complex, full-body activities characterized by multiple movement phases, rapid posture transitions, and simultaneous coordination of several limb groups. These dynamics introduce frequent occlusion, abrupt viewpoint changes, and more complex spatial relationships between joints, all of which challenge current VLM architectures. The pronounced drop in accuracy for these tasks underscores an important limitation: today’s VLMs are not yet adept at reliably parsing the fluid, multi-stage mechanics of compound movements.
Figure 4: Accuracy of VLM-generated Feedback Across Exercises: Top 3 vs. Bottom 3
5 The future of at-home fitness coaching
From Feedback to Collaboration: A key opportunity lies in shifting from one-way feedback to two-way collaboration. In its current form, FormCoach primarily tells the user what to fix, but true coaching is often a dialogue. We envision future AI fitness partners that not only give instructions, but also listen to the user’s concerns, goals, and progress [27, 28]. By integrating wearable sensing for real-time state detection with conversational AI, such systems could adapt feedback dynamically, offering not just corrections, but strategic guidance tailored to the user’s intent and physical condition.
Augmented Reality and Embodiment: Advances in AR hardware and software are rapidly opening the door to immersive, embodied coaching experiences [29, 30]. In an AR-enhanced FormCoach, lightweight smart glasses could project a virtual instructor directly into the user’s environment, visually overlaying highlights where their form diverges from the reference. Spatial audio could make feedback even more intuitive: imagine hearing your coach’s voice coming from the direction of a misaligned limb. This merging of spatial computing and AI analysis creates the potential for fully embodied, in-the-moment guidance that feels as natural as training with a human partner.
Accessible Fitness Coaching With AI coaches, expert guidance could become accessible to anyone with a camera, democratizing high-quality fitness training across geographies and income levels [31]. FormCoach envisions a future where every home gym includes a built-in AI trainer [32], much like AI assistants are now standard in smart homes [33, 34]. The challenge will be to ensure that as these systems become more pervasive, their feedback remains accurate, safe, and personalized—even as users push their physical limits.
6 Conclusion
We presented FormCoach, an interactive AI form coach capable of delivering targeted, actionable exercise feedback in real time. Alongside the system, we developed a curated dataset of expert-annotated user–reference exercise pairs and used it to systematically evaluate state-of-the-art vision–language models (VLMs) on nuanced, comparative human motion analysis. Our results demonstrate that VLMs can move beyond generic video understanding toward precise, context-aware guidance, positioning AI as an interactive, creative coaching partner. Our experiments show that while top-performing models such as GPT-4.1 exhibit strong actionability and partial accuracy, they still miss subtle form discrepancies and occasionally hallucinate nonexistent issues. These findings highlight both the promise and the limitations of current VLMs in high-stakes, precision-oriented applications. Overall, FormCoach opens the door to a new class of AI systems that not only see but also engage with users, guiding them through skill acquisition in a collaborative, human-like manner. As multimodal AI advances and incorporates richer sensing, the vision of universally accessible, trustworthy AI coaches moves closer to reality.
Appendix A Technical Appendices and Supplementary Material
A.1 Dataset
The dataset includes 22 common strength and mobility exercises as listed in Table 3. The dataset is hosted on Huggingface and can be accessed Here.
Table 3: List of Exercises Included in the Dataset
No. Exercise Name
1 Dumbbell Lift Overhead
2 Dumbbell Lunge Backward
3 Hands Up Rotate
4 Barbell Pull Up
5 Clean and Press
6 Dumbbell Lateral Raise
7 Burpee
8 Up and Down Stretch
9 Dumbbell Chest Fly
10 Arms Raise Overhead
11 Standing Ab Twists
12 Opposite Arms Up and Down Warmup
13 Dumbbell Squat and Press
14 Lunge and Reach Hands Front
15 Mule Kick
16 Arms Circles
17 Squat
18 One Arm Row
19 Squat and Jump with Hands Up
20 Dumbbell High Pulls
21 Deadlift
22 Barbell Shrug
A.2 Ground Truth Instruction Annotations
Experts annotated ground truth instructions given the user and reference exercise pairs through the Gradio interface as illustrated in Figure  5.
Figure 5: Instruction Annotation Interface on Gradio
A.3 Prompts
Prompts for feedback generation and evaluations are shown in Figure 6, Figure 8, Figure 8, and Figure 9.
Figure 6: Prompt for feedback generation
Figure 7: Prompt for accuracy evaluation
Figure 8: Prompt for actionability evaluation
Figure 9: Prompt for hallucination evaluation
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction describes the motivation, background, and main contributions of FormCoach which is an AI fitness coach powered by VLMs.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 4.1 and Section 5 discusses the limitation of video-only approach and future directions of improving form feedback accuracy.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: FormCoach provides the prompts for generating feedback and details about how the test dataset is generated and feedbacks are evaluated.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The gradio demo space is public for viewing the dataset. We’ll release a repo for the evaluation pipeline and the web demo soon.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The data can be accessed on the gradio app and the code can be accessed on github as mentioned in appendix.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The evaluation data, models and prompts are provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Evaluation metrics are thoroughly explained and results are discussed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section 4.2 mentions the resources used.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes] Yes
  44. Justification: Anonymity is preserved.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Broader impacts are discussed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper uses publicly open datasets and models and have no such concerns.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets used are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The dataset is hosted on huggingface.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: The annotator UI is provided in the appendix.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: IRB not required for our study on publicly available dataset.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Usage of LLM for evaluation is explained in the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.